# Coherent control of optomechanical entanglement and steering via dual parametric amplification

### 16 Apr 2026 v1

@@ Abstract
preprint: APS/123-QED
I Introduction:
Quantum entanglement [1, 2, 3, 4, 5], as one of the most fundamental features of quantum mechanics, is a key source for quantum metrology [6, 7], quantum computation [8, 9], and quantum communication[10, 11]. It has been demonstrated in atoms [12], ions [13], and cavity optomechanical (COM) systems [14, 15, 16, 17], where the interaction between an optical cavity and mechanical motion is mediated by radiation pressure [18, 19]. COM systems have enabled a variety of important phenomena and applications, such as mechanical squeezing [20, 21], optomechanically induced transparency [22, 23], and mechanical resonator cooling [24, 25]. Beyond entanglement, Einstein-Podolsky-Rosen (EPR) steering is a form of quantum nonlocality intermediate between entanglement and Bell nonlocality [26, 27]. Several schemes to generate steady-state one-way steering in COM systems have also been explored [28, 29]. To protect and enhance the fragile quantum correlations in COM platforms, substantial efforts have been devoted to developing various approaches. Among these, parametric driving enabled by optical and mechanical parametric amplification (OPA and MPA) provides a versatile and effective means to enhance nonlinearity and access rich classical and quantum dynamics [30, 31, 32, 33, 34, 35].
@@ I Introduction:
Inspired by pervious studies, we propose a scheme to enhance optomechanical entanglement and EPR steering in a COM system consisting of a nonlinear medium and a membrane embedded in a Fabry–Pérot (FP) cavity with a feedback loop. The loop is formed by a highly reflect mirror (HRM) and a controllable beam splitter (CBS), which feeds the cavity output field back into the cavity. The nonlinear medium and the membrane are driven by two laser fields to implement OPA and MPA, while an additional external laser field pumps the FP cavity through the CBS. Our proposal combines two effective ways and are expected to amplify quantum correlations. Within experimentally feasible parameter regimes, we first investigate the limit-cycle cavity dynamics and entanglement as functions of the CBS reflectivity. As the reflectivity increases, the effective cooperativity is enhanced and the system approaches the dynamical-instability threshold, leading to a larger cavity oscillation amplitude. Meanwhile, the maximum entanglement exhibits a non-monotonic dependence, increasing at first and then decreasing. This behavior indicates that a moderate reflectivity can be chosen to enable constructive interference of quantum-noise pathways between the output field and the input field, thereby amplifying the entanglement. We then demonstrate the joint modulation induced by parametric amplification and the coherent feedback loop. The maximum entanglement is evaluated as a function of the OPA-MPA modulation-frequency ratio (amplitude ratio) and reflectivity. Using the same set of parameters, we further evaluate the maximum steering and find coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables the transition between one-way and two-way steering in the COM system. By varying the frequency ratio, amplitude ratio, and reflectivity, one can enhance the steering and control its strength as well as the transition between one-way and two-way steering. Finally, we show the quantum correlations become more robust against thermal noise. The results provide physical insight into multi-field driven COM systems with coherent feedback. Our scheme offers an effective route to strong, high-quality, and widely tunable entanglement and steering, with potential applications in quantum computation and quantum communication.
II System and dynamics:
@@
Figure 1: Schematic diagram of the proposed COM system with the embedded membrane and the $\chi^{(2)}$ nonlinear medium, coupled to the coherent feedback loop. (a) The membrane and the nonlinear medium are each pumped by two laser fields with amplitude $G_{m}$ and $G_{c}$, frequency $\omega_{m}$ and $\omega_{c}$, respectively. (b)The coherent feedback loop consists of the HRM and the CBS. The external laser field with amplitude $E$ drives the cavity via the CBS. Here $a_{\mathrm{in}}$ denotes the optical input vacuum noise associated with the zero-point fluctuation entering the cavity through the CBS. The blue arrows indicate that the optical output field $a_{\mathrm{out}}$ reflected by the HRM is fed back into the cavity through the CBS.
@@
As shown in Fig. 1, we consider a COM setup in which a nonlinear $\chi^{(2)}$ medium and a membrane are embedded inside a Fabry–Pérot (FP) cavity with a feedback loop. The FP cavity supports a photon mode with frequency $\omega_{a}$. The membrane acting as a mechanical oscillator supports a phonon mode with effective mass $m$ and frequency $\omega_{b}$ [16]. The cavity is pumped by an external laser field with amplitude $E$, transmitted through a controllable beam splitter (CBS) with transmission $t_{b}$ and reflectivity $r_{b}$, satisfying $t_{b}^{2}+r_{b}^{2}=1$ [17]. $E=\sqrt{2\kappa_{a}P_{l}/(\hbar\omega_{l})}$ is the amplitude of the laser, where $\kappa_{a}$ is the decay rate of the optical cavity, $P_{l}$ ($\omega_{l}$) is the laser power (frequency) [47]. A coherent field with amplitude $G_{c}$, frequency $\omega_{c}$, and phase $\theta_{c}$ is applied to the nonlinear medium inside the cavity, giving rise to optical parametric amplification (OPA) [30, 31]. The membrane is simultaneously subjected to a periodic pump of frequency $\omega_{m}$, amplitude $G_{m}$, and phase $\theta_{m}$, which periodically modulates the spring constant and produces mechanical parametric amplification (MPA) [35, 15]. The Hamiltonian of the COM systems in a frame rotating with respect to the driving laser frequency $\omega_{l}$ reads [30]
@@
In the feedback channel, the output field is reflected by the HRM and then fed back into the optical cavity through the beam splitter. By using the standard input-output relation, the output field is given by $a_{\mathrm{out}}=\sqrt{2\kappa_{a}}a-a_{\mathrm{in}}$ [48]. The feedback-modified input field can be described as a superposition of the original input noise $a_{in}$ and the returned output field. These two contributions are coherently mixed by a lossless CBS before re-entering the cavity, so that the input-field operator can be written as: $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}=r_{b}e^{\mathrm{i}\theta}a_{\mathrm{out}}+t_{b}a_{\mathrm{in}}.$ $\theta$ denotes the phase shift acquired by the output field upon reflection from the HRM. By replacing the original input noise operator $a_{\mathrm{in}}$ by $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}$, the dynamical evolution of the COM system are given by a set of quantum Langevin equations(QLEs):
@@
where $\Gamma_{m}=2G_{m}\mathrm{cos}(\omega_{m}t-\theta_{m})$, $\Gamma_{a}=2G_{a}\mathrm{cos}(\Delta_{a}t-\theta_{a})$, $\zeta_{m}=2G_{m}\mathrm{sin}(\omega_{m}t-\theta_{m})$, $\zeta_{a}=2G_{a}\mathrm{sin}(\Delta_{a}t-\theta_{a})$. We can find the dynamics of the system can be effected by the effects of OPA, MPA, and the coherent feedback loop. Therefore, we show that the dynamics of the system can be regulated by the reflectivity of the CBS in Fig. 2. The dynamics of the quantum fluctuations can be fully characterized by the covariance matrix (CM) $\mathbf{V}$, a $4\times 4$ real symmetric matrix, with its matrix elements defined as $V_{ij}=\frac{1}{2}\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle$,($i,j=1,2,3,4$). The equation of motion of the CM ($V(t)$) is,
@@
When studying the squeezing feature of the phonon mode, the squeezing degree of a quadrature component can be characterized as (in units of $\mathrm{dB}$)
@@
In addition, the purity of the phonon mode is defined as
@@
III Numerical simulations and discussions:
In this section, we study how to control quantum correlations by the synergistic effect of OPA, MPA, and coherent feedback based on numerical simulations. We numerically study how coherent feedback and key system parameters affect quantum correlations in the COM system. We first analyze the bipartite entanglement, with particular emphasis on the roles of coherent feedback, the OPA-MPA modulation-frequency ratio, and the OPA-MPA amplitude ratio. We then analyze steering, focusing on its periodic dynamics, its dependence on the CBS reflectivity and on the OPA–MPA modulation-frequency and amplitude ratios. At least, we analyze the robustness of entanglement and steering against thermal noise. We choose experimentally feasible parameters within the range of typical values [15, 31]: $\omega_{b}/2\pi=1~\mathrm{MHz}$, $\kappa_{a}/2\pi=0.5~\mathrm{MHz}$, $\kappa_{b}/2\pi=1~\mathrm{Hz}$, $E/2\pi=60~\mathrm{GHz}$, $\lambda_{l}=1550~\mathrm{nm}$, $P_{l}\approx 2.9~\mathrm{mW},$ $m=150~\mathrm{ng}$, $g/2\pi=4~\mathrm{Hz}$ and $T=20~\mathrm{mK}$. In the process of numerical simulation, we keep all eigenvalues of matrix $A(t)$ with negative real parts during the numerical simulations.
@@ III Numerical simulations and discussions:
Figure 3: Maximum entanglement $E_{N,\mathrm{max}}$ as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (b). The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $G_{m}/G_{c}=1.5$ in (a) and $\omega_{m}/\Delta_{c}=1.6$ in (b). The other parameters are the same as those in Fig. 2.
@@
Figs. 3(a) and 3(b) demonstrate the combined modulation effects of the parametric amplifications and the coherent feedback loop. As shown in Fig. 3(a), the entanglement $E_{N,\mathrm{max}}$ exhibits a pronounced maximum when the frequency of the MPA satisfies $\omega_{m}\approx 1.7\Delta_{c}\approx 2\omega_{b}$. In this situation, the MPA term $\mathrm{i}G_{m}(b^{\dagger 2}e^{-\mathrm{i}\omega_{m}t}-b^{2}e^{\mathrm{i}\omega_{m}t})$ becomes time independent in the rotating frame of the phonon mode. This yields an effective degenerate squeezing Hamiltonian $H_{\mathrm{MPA}}^{\mathrm{eff}}=\mathrm{i}G_{m}(b^{\dagger 2}-b^{2})$, which maximize quadrature squeezing and suppress the thermal noise. The condition is a requirement for phase sensitive amplification in a harmonic oscillator and has been widely used in periodically modulated COM schemes to enhance quantum correlations. The effective decay rate decreases as the reflectivity increases, which has an impact on the entanglement of system. For a fixed modulation frequency ratio, it’s a feasible way to optimize the entanglement by tuning the reflectivity to an appropriate value. However, the entanglement is reduced when the reflectivity approaches $0.35$. Besides, it is observed that many stripes are also in Fig. 3(a). Fig. 3(b) illustrates the influence of the amplitude ratio and the reflectivity on the entanglement. In the absence of both coherent feedback and MPA, the entanglement is almost minimal. When we keep the amplitude of OPA fixed, the entanglement can reach a maximum of $0.42$ with the increase of amplitude of MPA. Besides, the impact of coherent feedback not only makes the tunable parametric area larger, but also enable the entanglement achieves a higher maximum value of $0.52$.
@@
Figure 4: Gaussian steering as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a) and (b), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (c) and (d). The white line denotes the boundary of zero steering. The parameters are $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $G_{m}=0.05\,\omega_{b}$ in (a) and (b), and $E/2\pi=60~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $\omega_{m}/\Delta_{c}=1.7$ in (c) and (d). The other parameters are the same as those in Fig. 3.
@@
In Fig. 4, we investigate the influence of the frequency ratio, amplitude ratio, and CBS reflectivity on the steering in the COM system. As shown in Fig. 4(a)-(b), only one-way steering can be achieved without the coherent feedback. The introduction of coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables a reflectivity-controlled transition between one-way and two-way steering in the COM system. To better show the transition, we also plot several slices at fixed $\omega_{m}/\Delta_{c}$. From these slices, one can see that $G_{b\to a,\max}$ remains finite over the whole range of $r_{b}$, while $G_{a\to b,\max}$ emerges only when $r_{b}$ exceeds a threshold value and then increases. Besides, the maximum steering appears near the area at $r_{b}=0.2$ and $\omega_{m}/\Delta_{c}=0.9$$1.3$, where the parameter is close the stability threshold. This behavior can be understood as follows. The reflectivity directly modifies the effective decay rate of the photon mode and its corresponding matrix elements in the diffusion matrix, while the matrix elements of the phonon mode remain unchanged. There is another obvious phenomenon in which the steering suddenly drop to zero as the $\omega_{m}/\Delta_{c}$ approaches $1$. As we mentioned before, this is because the steering criterion is no longer satisfied due to a sudden change which enables the $\det V_{a}/\det V_{ab}<1$ around this parameter region. Similar abrupt behavior also appears in the numerical calculation of entanglement. Figs. 4(c)-(d) show the steering as functions of the CBS reflectivity and the amplitude ratio. It should be noted that the reflectivity is extended up to 0.3 because the stability condition is less restrictive. By varying the ratio $G_{m}/G_{c}$, one can also enhance the steering and achieve the transition between one-way and two-way steering. Overall, Fig. 4 reveals the combined modulation effects of the frequency ratio, the amplitude ratio, and coherent feedback on steering in the COM system. The CBS reflectivity determines asymmetric steering resulted from the asymmetry of the system [52]. These results show that coherent feedback not only enlarges the steering region, but also enables control over the steering strength and type through the tuning of system parameters. In particular, the existence of one-way steering in our proposed COM system may have promising applications in one-sided device-independent quantum key distribution [53, 54].
@@
Figure 5: Purity and squeezing of the phonon mode as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$. Panel (a) shows the purity, and panel (b) shows the squeezing. The parameters are $E/2\pi=50~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, and $G_{m}/G_{c}=1.5$ in both panels. The other parameters are the same as those in Fig. 4.
@@
High purity is a prerequisite for many mechanical quantum effects. For example, back-action-evading measurements or squeezing require the oscillator to be prepared in a high-purity quantum state [53, 54]. Therefore, in order to show the local quantum characteristics of the phonon mode, we also investigate how the purity and squeezing change with the reflectivity and the frequency ratio in Fig. 5. From Fig. 5(a), the purity sustains a relatively high level over almost the whole figure and only shows a small area with low value around $\omega_{m}/\Delta_{c}=1.7$. With coherent feedback, it’s useful to adjust the reflectivity so as to turn the original low value area into high value area. Fig. 5(b) shows a nearly opposite phenomenon, where most of the area has relatively low values, but the squeezing is almost the highest at $\omega_{m}/\Delta_{c}=1.7$ without coherent feedback. This can also be understood for the similar reason discussed as before. When $\omega_{m}/\Delta_{c}\approx 1.7$, $\omega_{m}$ becomes close to $2\omega_{b}$, which makes the enhancement of MPA induced squeezing on the phonon mode more pronounced. In the other high squeezing area, we can adjust the reflectivity to broaden this region. Compared with Fig. 5(a) and Fig. 5(b), we can find that, by introducing coherent feedback and properly tuning the parameters, the local quantum characteristics of the phonon mode can maintain relatively high purity while still exhibiting strong squeezing. This indicates that the present scheme provides favorable initial-state conditions for observing quantum behavior of the COM systems and offers a promising platform for the preparation of non-classical states of motion, with potential applications in quantum metrology and precision sensing [53, 54, 55].
@@
Figure 6: Maximum entanglement and steering as functions of the feedback phase $\theta$ and the modulation-amplitude ratio $G_{m}/G_{c}$. Panels (a), (b) show $E_{N,\max}$ and $G_{a\rightarrow b,\max}$ respectively. The parameters are $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.2$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.15$, and $\omega_{m}/\Delta_{c}=1.3$ in (b). The parameters are the same as those in Fig. 5.
@@
Figure 7: Thermal robustness of entanglement and steering under coherent feedback. Panels (a), (b), and (c) show $E_{N,\max}$, $G_{a\rightarrow b,\max}$, and $G_{b\rightarrow a,\max}$, respectively, as functions of the environment temperature $T$ and the reflectivity $r_{b}$. The parameters are $E/2\pi=50~\mathrm{GHz}$, $\Delta_{c}=1.18\,\omega_{b}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=1.5$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=3.0$, and $\omega_{m}/\Delta_{c}=1.5$ in (b) and (c). The other parameters are the same Fig. 6.

### 18 Apr 2026 v2

@@ Abstract
keywords:
optomechanics, quantum correlations, coherent feedback, parametric amplification
articletype: Paper
1 Introduction:
Quantum entanglement [1, 2, 3, 4, 5], as one of the most fundamental features of quantum mechanics, is a key source for quantum metrology [6, 7], quantum computation [8, 9], and quantum communication[10, 11]. It has been demonstrated in atoms [12], ions [13], and cavity optomechanical (COM) systems [14, 15, 16, 17, 18], where the interaction between an optical cavity and mechanical motion is mediated by radiation pressure [19, 20]. COM systems have enabled a variety of important phenomena and applications, such as mechanical squeezing [21, 22], optomechanically induced transparency [23, 24], and mechanical resonator cooling [25, 26]. Beyond entanglement, Einstein-Podolsky-Rosen (EPR) steering is a form of quantum nonlocality intermediate between entanglement and Bell nonlocality [27, 28]. Several schemes to generate steady-state one-way steering in COM systems have also been explored [29, 30]. To protect and enhance the fragile quantum correlations in COM platforms, substantial efforts have been devoted to developing various approaches. Among these, parametric driving enabled by optical and mechanical parametric amplification (OPA and MPA) provides a versatile and effective means to enhance nonlinearity and access rich classical and quantum dynamics [31, 32, 33, 34, 35, 36].
@@ 1 Introduction:
Inspired by previous studies, we propose a scheme to control optomechanical entanglement and EPR steering in a COM system consisting of a nonlinear medium and a membrane embedded in a Fabry–Pérot (FP) cavity with a feedback loop. The loop is formed by a highly reflect mirror (HRM) and a controllable beam splitter (CBS), which feeds the cavity output field back into the cavity. The nonlinear medium and the membrane are driven by two laser fields to implement OPA and MPA, while an additional external laser field pumps the FP cavity through the CBS. Our proposal combines two effective ways and are expected to amplify quantum correlations. Within experimentally feasible parameter regimes, we first investigate the limit-cycle cavity dynamics and entanglement as functions of the CBS reflectivity. As the reflectivity increases, the effective cooperativity is enhanced and the system approaches the dynamical-instability threshold, leading to a larger cavity oscillation amplitude. Meanwhile, the maximum entanglement exhibits a non-monotonic dependence, increasing at first and then decreasing. This behavior indicates that a moderate reflectivity can be chosen to enable constructive interference of quantum-noise pathways between the output field and the input field, thereby amplifying the entanglement. We then demonstrate the joint modulation induced by parametric amplification and the coherent feedback loop. The maximum entanglement is evaluated as a function of the OPA-MPA modulation-frequency ratio (amplitude ratio) and reflectivity. Using the same set of parameters, we further evaluate the maximum steering and find coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables the transition between one-way and two-way steering in the COM system. By varying the frequency ratio, amplitude ratio, and reflectivity, one can enhance the steering and control its strength as well as the transition between one-way and two-way steering. Finally, we show the quantum correlations become more robust against thermal noise. The results provide physical insight into multi-field driven COM systems with coherent feedback. Our scheme offers an effective route to strong, high-quality, and widely tunable entanglement and steering, with potential applications in quantum computation and quantum communication.
2 System and dynamics:
@@
Figure 1: Schematic diagram of the proposed COM system with the embedded membrane and the $\chi^{(2)}$ nonlinear medium, coupled to the coherent feedback loop. (a) The membrane and the nonlinear medium are each pumped by two laser fields with amplitudes $G_{m}$ and $G_{c}$, frequencies $\omega_{m}$ and $\omega_{c}$, respectively. (b) The coherent feedback loop consists of the HRM and the CBS. The external laser field with amplitude $E$ drives the cavity via the CBS. Here $a_{\mathrm{in}}$ denotes the optical input vacuum noise associated with the zero-point fluctuation entering the cavity through the CBS. The blue arrows indicate that the optical output field $a_{\mathrm{out}}$ reflected by the HRM is fed back into the cavity through the CBS.
@@
As shown in Fig. 1, we consider a COM setup in which a nonlinear $\chi^{(2)}$ medium and a membrane are embedded inside a Fabry–Pérot (FP) cavity with a feedback loop. The FP cavity supports a photon mode with frequency $\omega_{a}$. The membrane acting as a mechanical oscillator supports a phonon mode with effective mass $m$ and frequency $\omega_{b}$ [16]. The cavity is pumped by an external laser field with amplitude $E$, transmitted through a controllable beam splitter (CBS) with transmission $t_{b}$ and reflectivity $r_{b}$, satisfying $t_{b}^{2}+r_{b}^{2}=1$ [17]. $E=\sqrt{2\kappa_{a}P_{l}/(\hbar\omega_{l})}$ is the amplitude of the laser, where $\kappa_{a}$ is the decay rate of the optical cavity, $P_{l}$ ($\omega_{l}$) is the laser power (frequency) [48]. A coherent field with amplitude $G_{c}$, frequency $\omega_{c}$, and phase $\theta_{c}$ is applied to the nonlinear medium inside the cavity, giving rise to optical parametric amplification (OPA) [31, 32]. The membrane is simultaneously subjected to a periodic pump of frequency $\omega_{m}$, amplitude $G_{m}$, and phase $\theta_{m}$, which periodically modulates the spring constant and produces mechanical parametric amplification (MPA) [36, 15]. The Hamiltonian of the COM system in a frame rotating with respect to the driving laser frequency $\omega_{l}$ reads [31]
@@
In the feedback channel, the output field is reflected by the HRM and then fed back into the optical cavity through the beam splitter. By using the standard input-output relation, the output field is given by $a_{\mathrm{out}}=\sqrt{2\kappa_{a}}a-a_{\mathrm{in}}$ [49]. The feedback-modified input field can be described as a superposition of the original input noise $a_{in}$ and the returned output field. These two contributions are coherently mixed by a lossless CBS before re-entering the cavity, so that the input-field operator can be written as: $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}=r_{b}e^{\mathrm{i}\theta}a_{\mathrm{out}}+t_{b}a_{\mathrm{in}}.$ $\theta$ denotes the phase shift acquired by the output field upon reflection from the HRM. By replacing the original input noise operator $a_{\mathrm{in}}$ by $\tilde{a}_{\mathrm{fb}}^{\mathrm{in}}$, the dynamical evolution of the COM system is given by a set of quantum Langevin equations (QLEs):
@@
where $\Gamma_{m}=2G_{m}\mathrm{cos}(\omega_{m}t-\theta_{m})$, $\Gamma_{a}=2G_{a}\mathrm{cos}(\Delta_{a}t-\theta_{a})$, $\zeta_{m}=2G_{m}\mathrm{sin}(\omega_{m}t-\theta_{m})$, $\zeta_{a}=2G_{a}\mathrm{sin}(\Delta_{a}t-\theta_{a})$. Equation 5 the dynamics of the system can be affected by the effects of OPA, MPA, and the coherent feedback loop. Therefore, we show that the dynamics of the system can be regulated by the reflectivity of the CBS in Fig. 2. The dynamics of the quantum fluctuations can be fully characterized by the covariance matrix (CM) $\mathbf{V}$, a $4\times 4$ real symmetric matrix, with its matrix elements defined as $V_{ij}=\frac{1}{2}\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle$,($i,j=1,2,3,4$). The equation of motion of the CM ($V(t)$) is,
@@
When studying the squeezing feature of the phonon mode, the squeezing degree of a quadrature component can be characterized as [51] (in units of $\mathrm{dB}$)
@@
In addition, the purity of the phonon mode is defined as [52]
@@
3 Numerical simulations and discussions
In this section, we study how to control quantum correlations by the synergistic effect of the OPA, MPA, and coherent feedback based on numerical simulations. We numerically study how coherent feedback and the OPA-MPA modulation frequency (amplitude) ratio affect quantum correlations in the COM system. Moreover, we study the local mechanical quantum property of the phonon mode. At last, we analyze the modulation of the phase shift and the robustness of entanglement and steering against thermal noise. We choose experimentally feasible parameters within the range of typical values [15, 32]: $\omega_{b}/2\pi=1~\mathrm{MHz}$, $\kappa_{a}/2\pi=0.5~\mathrm{MHz}$, $\kappa_{b}/2\pi=1~\mathrm{Hz}$, $E/2\pi=60~\mathrm{GHz}$, $\lambda_{l}=1550~\mathrm{nm}$, $P_{l}\approx 2.9~\mathrm{mW},$ $m=150~\mathrm{ng}$, $g/2\pi=4~\mathrm{Hz}$ and $T=20~\mathrm{mK}$. In the process of numerical simulation, we keep all eigenvalues of matrix $A(t)$ with negative real parts during the numerical simulations.
@@ 3 Numerical simulations and discussions
Figure 3: Maximum entanglement $E_{N,\mathrm{max}}$ as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (b). Here, $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $G_{m}/G_{c}=1.5$ in (a) and $\omega_{m}/\Delta_{c}=1.6$ in (b). The other parameters are the same as those in Fig. 2.
@@
Figs. 3(a) and 3(b) demonstrate the combined modulation effects of the parametric amplifications and the coherent feedback loop. As shown in Fig. 3(a), the entanglement $E_{N,\mathrm{max}}$ exhibits a pronounced maximum when the frequency of the MPA satisfies $\omega_{m}\approx 1.7\Delta_{c}\approx 2\omega_{b}$. In this situation, the MPA term $\mathrm{i}G_{m}(b^{\dagger 2}e^{-\mathrm{i}\omega_{m}t}-b^{2}e^{\mathrm{i}\omega_{m}t})$ becomes time independent in the rotating frame of the phonon mode. This yields an effective degenerate squeezing Hamiltonian $H_{\mathrm{MPA}}^{\mathrm{eff}}=\mathrm{i}G_{m}(b^{\dagger 2}-b^{2})$, which maximizes quadrature squeezing and suppress the thermal noise [55]. The condition is a requirement for phase sensitive amplification in a harmonic oscillator and has been widely used in periodically modulated COM schemes to enhance quantum correlations. The effective decay rate decreases as the reflectivity increases, which has an impact on the entanglement of system. For a fixed modulation frequency ratio, it’s a feasible way to optimize the entanglement by tuning the reflectivity to an appropriate value. However, the entanglement is reduced when the reflectivity approaches $0.35$. Fig. 3(b) illustrates the influence of the amplitude ratio and the reflectivity on the entanglement. In the absence of both coherent feedback and MPA, the entanglement is almost minimal. When we keep the amplitude of OPA fixed, the entanglement can reach a maximum of $0.42$ with the increase of amplitude of MPA. Besides, the impact of coherent feedback not only enlarges the tunable parametric region, but also enables the entanglement achieves a higher maximum value of $0.52$.
@@
Figure 4: Gaussian steering as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$ in (a) and (b), and as functions of $G_{m}/G_{c}$ and $r_{b}$ in (c) and (d). The white line denotes the boundary of zero steering. Here, $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $G_{m}=0.05\,\omega_{b}$ in (a) and (b), and $E/2\pi=60~\mathrm{GHz}$, $G_{c}=0.03\,\omega_{b}$, and $\omega_{m}/\Delta_{c}=1.7$ in (c) and (d). The other parameters are the same as those in Fig. 3.
@@
In Fig. 4, we investigate the influence of the frequency ratio, amplitude ratio, and CBS reflectivity on the steering in the COM system. As shown in Fig. 4(a)-(b), only one-way steering can be achieved without the coherent feedback. The introduction of coherent feedback not only enlarges the parameter regime in which one-way steering exists, but also enables a reflectivity-controlled transition between one-way and two-way steering in the COM system. To better show the transition, we also plot several slices at fixed $\omega_{m}/\Delta_{c}$. From these slices, one can see that $G_{b\to a,\max}$ remains finite over the whole range of $r_{b}$, while $G_{a\to b,\max}$ emerges only when $r_{b}$ exceeds a threshold value and then increases. Besides, the maximum steering appears near the area at $r_{b}=0.2$ and $\omega_{m}/\Delta_{c}=0.9$$1.3$, where the parameter is close to the stability threshold. This behavior can be understood as follows. The reflectivity directly modifies the effective decay rate of the photon mode and its corresponding matrix elements in the diffusion matrix, while the matrix elements of the phonon mode remain unchanged. There is another obvious phenomenon in which the steering suddenly drops to zero as the $\omega_{m}/\Delta_{c}$ approaches $1$. As we mentioned before, this is because the steering criterion is no longer satisfied due to a sudden change that makes the $\det V_{a}/\det V_{ab}<1$ around this parameter region. Similar abrupt behavior also appears in the numerical calculation of entanglement. Figs. 4(c)-(d) show the steering as functions of the CBS reflectivity and the amplitude ratio. It should be noted that the reflectivity is extended up to 0.3 because the stability condition is less restrictive. By varying the ratio $G_{m}/G_{c}$, one can also enhance the steering and achieve the transition between one-way and two-way steering. Overall, Fig. 4 reveals the combined modulation effects of the frequency ratio, the amplitude ratio, and coherent feedback on steering in the COM system. The CBS reflectivity determines asymmetric steering resulting from the asymmetry of the system [56]. These results show that coherent feedback not only enlarges the steering region, but also enables control over the steering strength and type through the tuning of system parameters. In particular, the existence of one-way steering in our proposed COM system may have promising applications in one-sided device-independent quantum key distribution [57, 58].
@@
Figure 5: Purity and squeezing of the phonon mode as functions of $\omega_{m}/\Delta_{c}$ and $r_{b}$. Panel (a) shows the purity, and panel (b) shows the squeezing. Here, $E/2\pi=50~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, and $G_{m}/G_{c}=1.5$ in both panels. The other parameters are the same as those in Fig. 4.
@@
High purity is a prerequisite for many mechanical quantum effects. For example, back-action-evading measurements or squeezing require the oscillator to be prepared in a high-purity quantum state [57, 58]. Therefore, in order to show the local quantum characteristics of the phonon mode, we also investigate how the purity and squeezing change with the reflectivity and the frequency ratio in Fig. 5. From Fig. 5(a), the purity remains at a relatively high level over almost the entire parameter region and only shows a small area with low value around $\omega_{m}/\Delta_{c}=1.7$. With coherent feedback, the reflectivity can be tuned so as to turn the original low value area into high value area. Fig. 5(b) shows a nearly opposite phenomenon, where most of the area has relatively low values, but the squeezing is almost the highest at $\omega_{m}/\Delta_{c}=1.7$ without coherent feedback. This can also be understood for the similar reason discussed above. When $\omega_{m}/\Delta_{c}\approx 1.7$, $\omega_{m}$ becomes close to $2\omega_{b}$, which makes the enhancement of MPA-induced squeezing on the phonon mode more pronounced. In the other high squeezing area, we can adjust the reflectivity to broaden this region. By comparing Fig. 5(a) and Fig. 5(b), we can find that, by introducing coherent feedback and properly tuning the parameters, the local quantum characteristics of the phonon mode can maintain relatively high purity while still exhibiting strong squeezing. This indicates that the present scheme provides favorable initial-state conditions for observing quantum behavior of the COM systems and offers a promising platform for the preparation of non-classical states of motion, with potential applications in quantum metrology and precision sensing [57, 58, 59].
@@
Figure 6: Maximum entanglement and steering as functions of the phase shift $\theta$ and the amplitude ratio $G_{m}/G_{c}$. Panels (a), (b) show $E_{N,\max}$ and $G_{a\rightarrow b,\max}$ respectively. Here, $E/2\pi=50~\mathrm{GHz}$ and $G_{c}=0.02\,\omega_{b}$ in both panels, with $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.2$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $\Delta_{c}=1.18\,\omega_{b}$, $r_{b}=0.15$, and $\omega_{m}/\Delta_{c}=1.3$ in (b). The parameters are the same as those in Fig. 5.
@@
Figure 7: Thermal robustness of entanglement and steering under coherent feedback. Panels (a), (b), and (c) show $E_{N,\max}$, $G_{a\rightarrow b,\max}$, and $G_{b\rightarrow a,\max}$, respectively, as functions of the environment temperature $T$ and the reflectivity $r_{b}$. Here, $E/2\pi=50~\mathrm{GHz}$, $\Delta_{c}=1.18\,\omega_{b}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=1.5$, and $\omega_{m}/\Delta_{c}=1.5$ in (a), and $E/2\pi=70~\mathrm{GHz}$, $G_{c}=0.02\,\omega_{b}$, $G_{m}/G_{c}=3.0$, and $\omega_{m}/\Delta_{c}=1.5$ in (b) and (c). The other parameters are the same Fig. 6.